Rocket-powered sleds are used to test the human response to acceleration. If a rocket-powered sled is acceleration to a speed of 444 m/s in 1.8 seconds, then what is the acceleration and what is the distance which the sled travels?
Answer by Alan3354(21605)

If a rocket-powered sled is acceleration to a speed of 444 m/s in 1.8 seconds, then what is the acceleration and what is the distance which the sled travels?
--------------------
a = v/t = 444m/s/1.8s
a = 2220/9 m/sec/sec
a = 246.666... m/sec^2
----------------------
s = (at^2)/2 = 246.666*1.8^2/2
s = 399.6 meters (distance traveled in 1.8 seconds)